# DEGAS: Detailed Expressions on Full-Body Gaussian Avatars

### v1

@@ Abstract
Although neural rendering has made significant advancements in creating lifelike, animatable full-body and head avatars, incorporating detailed expressions into full-body avatars remains largely unexplored. We present DEGAS, the first 3D Gaussian Splatting (3DGS)-based modeling method for full-body avatars with rich facial expressions. Trained on multiview videos of a given subject, our method learns a conditional variational autoencoder that takes both the body motion and facial expression as driving signals to generate Gaussian maps in the UV layout. To drive the facial expressions, instead of the commonly used 3D Morphable Models (3DMMs) in 3D head avatars, we propose to adopt the expression latent space trained solely on 2D portrait images, bridging the gap between 2D talking faces and 3D avatars. Leveraging the rendering capability of 3DGS and the rich expressiveness of the expression latent space, the learned avatars can be reenacted to reproduce photorealistic rendering images with subtle and accurate facial expressions. Experiments on an existing dataset and our newly proposed dataset of full-body talking avatars demonstrate the efficacy of our method. We also propose an audio-driven extension of our method with the help of 2D talking faces, opening new possibilities to interactive AI agents. Project page: https://initialneil.github.io/DEGAS .
@@ 1 Introduction
A naive way to enable facial expressions on full-body avatars would be using 3DMMs for facial control, for example, controlling the expression parameters of SMPL-X [45], or using the parameters of FLAME [35] or BFM [46] as driving signals. We find two drawbacks of such naive methods. 1) The expressiveness of 3DMMs is limited [43, 64, 58]. Being coarse meshes, 3DMMs are essentially not able to capture subtle facial changes. 2) Driving signal generation for a 3DMM is non-trivial. Both video-driven [77, 16, 40] and audio-driven [15, 64, 57, 49, 1] methods suffer from efficiency, accuracy, and expressiveness issues.
@@
Though both trained with registered meshes, head avatars and full-body avatars usually face very different registration quality. With well aligned mesh surfaces, the binding of the 3D Gaussians and the underlying mesh is usually simple for head avatars. Both SplattingAvatar [55] and GaussianAvatars [51] propose to bind 3D Gaussians to FLAME mesh triangles without any pose-dependent compensations. Full-body avatars, on the other hand, usually face the challenge of a much less accurate underlying mesh because of clothing. D3GA [78] proposes to alleviate this problem by modeling clothes with separate tetrahedron layers. We follow the practice of AnimatableGaussians [37] and CodecAvatars [2] to leverage the ability of 2D CNNs for the pose-dependent generation of 3DGS parameters.
@@
  • We propose driving 3D avatars with 2D talking faces, bridging the gap between these two research topics and opening new possibilities for the reenactment of photorealistic avatars.
@@ 2.3 Choice of Driving Signal
Different from a 4D playback system like FVV [10], or 4DGS [63], avatar modeling methods focus on the reenactment from accessible driving signals. For full-body avatars, apart from the commonly used skeleton poses as driving signal [52, 28, 24], AnimatableGaussians [37] also encodes the posed position maps. SurMo [26] further considers temporal dynamics to overcome the limitation of the deterministic nature of driving signals. Head avatars [79, 74, 51, 55], on the other hand, usually take into account the expression parameters of the underlying 3DMMs for better capturing the surface deformation of the face mesh.
@@ 3.1 Preliminaries: 3D Gaussian Splatting
where the rotation matrix $R_{i}$ and scaling matrix $S_{i}$ that formulate the covariance matrix $\Sigma_{i}$ are constructed from rotation quaternion $\boldsymbol{q}_{i}$ and scaling vector $\boldsymbol{s}_{i}$ respectively.
@@ 3.2 Driving Signal
Body signal. We use the body pose parameters from SMPL-X as the body signal which enables body motion control down to the finger level. We keep the motion of the head to make the avatar more natural. The jaw pose, on the other hand, is neutralized, leaving full facial control to the face signal.
@@ 3.3 Conditional Variational Autoencoder
Expression Encoder. DPE [43] proposes a bidirectional cyclic training strategy in the training process aiming at the disentanglement of pose and expression. We take the pretrained expression encoder as our facial controller. From the input multiview images, we apply the head pose estimator from SynergyNet [62] to find front-facing views. In the training stage, we choose the four most frontal views to extract expression codes. For every iteration, a randomly weighted average of the four codes is used as a face signal, as in Figure 4.
@@ 3.5 Loss and Training
Figure 8: Comparison on our DREAMS-Avatar dataset.
@@ 4 Experiments
ActorsHQ Dataset. ActorsHQ is a high quality dataset for full-body avatars. We follow the experiment setup of AnimatableGaussians [37] to use 47 full-body views (46 views for training and 1 for testing, 4x down-sampled images).
@@ 4.1 Comparison on Full-body Avatars
Figure 7 shows the qualitative comparison on ActorsHQ. The rendering results of our method exhibit rich texture details. The quantitative comparison is reported in Table 1. We report PSNR and SSIM calculated in the rendered full images, LPIPS [72] and FID [22] in the cropped regions. Our method quantitatively outperforms all three SoTA methods. One key observation we make is that the adoption of powerful 2D CNN networks in both AnimatableGaussians [37] and ours significantly improves pose-dependent modeling. Please see more discussions in Section 4.3.
@@ 4.2 Comparison on DREAMS-Avatar
View Synthesis. Table 2 shows the quantitative results of view synthesis on Sequence1 Cam02’s 500–1000 frames where both large poses and rich facial expressions are performed. We show the qualitative comparison in Figure 8. AnimatableGaussians [37] can render a high-quality body and a neutral face. PoseVocab [36] and GaussianAvatar [24] fail to model the face region due to the large expression variations in the dataset. While our method renders high-quality body and face details. We show more rendering results of DEGAS in Figure 6.
@@
Both our method and AnimatableGaussians [37] learn from CodecAvatars [2] in adopting powerful 2D CNN networks for the Gaussian parameters generation. The difference is that ours does not use skip connections between encoder and decoder, considering the potential steaming scenario [39]. Also, the bottleneck between the encoder and decoder that follows the UV layout of SMPL-X enables the face signal injection in our method. The runtime is compared in Table 3 where our method reaches a real-time framerate.
@@ 4.4 Ablation Study
Table 4: Choice of face signal. The 2D talking faces-based expression encoder from DPE [43] is a better choice as the face signal.
@@
Choice of face signal. We conducted ablation experiments to validate the design choice of using a pretrained expression encoder from DPE [43] as the face signal. We compared to using a 3DMM as the face signal by extracting FLAME parameters from the frontal camera. We replace the jaw pose of SMPL-X with that from the extracted FLAME parameters and use the expression parameters to condition our convolutional decoder. Note that in our method, the jaw pose is neutralized, i.e., the jaw movement is solely represented by the cVAE instead of mesh movements.
@@
The use of a pretrained encoder from 2D talking faces lets our method inherit the ability to both video-driven and audio-driven reenactment. With the help of SadTalker [73], we show the extension of DEGAS to full-body talking avatars. Given an audio clip, we firstly use SadTalker [73] to generate 2D talking videos from one face image of the subject, and then use the DPE [43] expression encoder to extract face signal to drive our avatars. The body pose generation is whole another research track [68, 42, 9]. Here we only showcase the generation from TalkSHOW [68] in Figure 11. Noted that the jaw pose generated by TalkSHOW is not used.
@@ 5 Conclusion
In this paper, we proposed DEGAS, the first 3DGS-based method for full-body avatars with subtle and accurate facial expressions. We discussed in this paper that an expression latent space pretrained solely on 2D talking faces, is a better choice for the reenactment of 3D avatars, opening new possibilities to interactive life-like agents. The avatars modeled with DEGAS can be animated and rendered in real-time framerate to perform natural body motions and rich facial expressions. We conducted qualitative and quantitative experiments to validate the efficacy of our method. We also showed the audio-driven extension of our method to demonstrate its potential for downstream applications.
@@ C Ethics
We captured six human subjects in our proposed DREAMS-Avatar dataset. All subjects have given written consents for the captured footage in this dataset. We make the data publicly available for research purposes.

### v2

@@ Abstract
Although neural rendering has made significant advances in creating lifelike, animatable full-body and head avatars, incorporating detailed expressions into full-body avatars remains largely unexplored. We present DEGAS, the first 3D Gaussian Splatting (3DGS)-based modeling method for full-body avatars with rich facial expressions. Trained on multiview videos of a given subject, our method learns a conditional variational autoencoder that takes both the body motion and facial expression as driving signals to generate Gaussian maps in the UV layout. To drive the facial expressions, instead of the commonly used 3D Morphable Models (3DMMs) in 3D head avatars, we propose to adopt the expression latent space trained solely on 2D portrait images, bridging the gap between 2D talking faces and 3D avatars. Leveraging the rendering capability of 3DGS and the rich expressiveness of the expression latent space, the learned avatars can be reenacted to reproduce photorealistic rendering images with subtle and accurate facial expressions. Experiments on an existing dataset and our newly proposed dataset of full-body talking avatars demonstrate the efficacy of our method. We also propose an audio-driven extension of our method with the help of 2D talking faces, opening new possibilities for interactive AI agents. Project page: https://initialneil.github.io/DEGAS .
@@ 1 Introduction
A naive way to enable facial expressions on full-body avatars would be using 3DMMs for facial control, for example, controlling the expression parameters of SMPL-X [46], or using the parameters of FLAME [36] or BFM [47] as driving signals. We find two drawbacks of such naive methods. 1) The expressiveness of 3DMMs is limited [44, 65, 59]. Being coarse meshes, 3DMMs are essentially not able to capture subtle facial changes. 2) Driving signal generation for a 3DMM is non-trivial. Both video-driven [78, 17, 41] and audio-driven [16, 65, 58, 50, 1] methods suffer from efficiency, accuracy, and expressiveness issues. Recent works [52, 43] have achieved photorealistic facial modeling in a very dense multi-view camera set-up. However, accurate 3DMM registration is non-trivial from a monocular camera [17, 41].
@@
Though both trained with registered meshes, head avatars and full-body avatars usually face very different registration qualities. With well aligned mesh surfaces, the binding of the 3D Gaussians and the underlying mesh is usually simple for head avatars. Both SplattingAvatar [56] and GaussianAvatars [52] propose to bind 3D Gaussians to FLAME mesh triangles without any pose-dependent compensations. Full-body avatars, on the other hand, usually face the challenge of a much less accurate underlying mesh because of clothing. D3GA [79] proposes to alleviate this problem by modeling clothes with separate tetrahedron layers. We follow the practice of AnimatableGaussians [38] and CodecAvatars [2] to leverage the ability of 2D CNNs for the pose-dependent generation of 3DGS parameters.
@@
  • We propose to drive 3D avatars with 2D talking faces, bridging the gap between these two research topics and opening new possibilities for the reenactment of photorealistic avatars.
@@ 2.3 Choice of Driving Signal
Different from a 4D playback system like FVV [10], or 4DGS [64], avatar modeling methods focus on the reenactment from accessible driving signals. For full-body avatars, apart from the commonly used skeleton poses as driving signals [53, 29, 25], AnimatableGaussians [38] also encodes the posed position maps. SurMo [27] further considers temporal dynamics to overcome the limitation of the deterministic nature of driving signals. Head avatars [80, 75, 52, 56], on the other hand, usually take into account the expression parameters of the underlying 3DMMs for better capturing the surface deformation of the face mesh. Audio2Photoreal [43] models mesh-based full-body avatars with facial control of accurate expression parameters registered from a dense camera set-up, which is non-trivial to acquire from a more casual set-up.
@@ 3.1 Preliminaries: 3D Gaussian Splatting
where the rotation matrix $R_{i}$ and scaling matrix $S_{i}$ that formulate the covariance matrix $\Sigma_{i}$ are constructed from the rotation quaternion $\boldsymbol{q}_{i}$ and the scaling vector $\boldsymbol{s}_{i}$ respectively.
@@ 3.2 Driving Signal
Body signal. We use the body pose parameters from SMPL-X as the body signal, which enables body motion control down to the finger level. We keep the motion of the head to make the avatar more natural. The jaw pose, on the other hand, is neutralized, leaving full facial control to the face signal.
@@ 3.3 Conditional Variational Autoencoder
Expression Encoder. DPE [44] proposes a bidirectional cyclic training strategy in the training process aiming at the disentanglement of pose and expression. We take the pretrained expression encoder as our facial controller. From the input multiview images, we apply the head pose estimator from SynergyNet [63] to find front-facing views. In the training stage, we choose the four most frontal views to extract expression codes. For every iteration, a randomly weighted average of the four codes is used as the face signal, as illustrated in Figure 4.
@@ 3.5 Loss and Training
Figure 8: Comparison on our DREAMS-Avatar dataset. Our method renders high quality details comparing to SoTA methods.
@@ 4 Experiments
ActorsHQ Dataset. ActorsHQ is a high-quality dataset for full-body avatars. We follow the experiment setup of AnimatableGaussians [38] to use 47 full-body views (46 views for training and 1 for testing, at 1k resolution).
@@ 4.1 Comparison on Full-body Avatars
Figure 7 shows the qualitative comparison on ActorsHQ. The rendering results of our method exhibit rich texture details. The quantitative comparison is reported in Table 1. We report PSNR and SSIM calculated in the rendered full images, LPIPS [73] and FID [23] in the cropped regions. Our method quantitatively outperforms all three SoTA methods. One key observation we make is that the adoption of powerful 2D CNN networks in both AnimatableGaussians [38] and ours significantly improves pose-dependent modeling. More discussions are in Section 4.3.
@@ 4.2 Comparison on DREAMS-Avatar
View Synthesis. Table 2 shows the quantitative results of view synthesis on Sequence1 Cam02’s 500–1000 frames where both large poses and rich facial expressions are performed. The facial expression accuracy is evaluated by the Average Expression Distance (AED) proposed in PIRenderer [54], which estimates the cosine distance of the expression coefficients extracted by Deep3DFaceRecon [12]. We show the qualitative comparison in Figure 8. AnimatableGaussians [38] can render a high-quality body and a neutral face. PoseVocab [37] and GaussianAvatar [25] fail to model the face region due to the large expression variations in the dataset. While our method renders high-quality body and face details. We show more rendering results of DEGAS in Figure 6 with various poses and rich facial expressions.
@@
Both our method and AnimatableGaussians [38] learn from CodecAvatars [2] in adopting powerful 2D CNN networks for the generation of Gaussian maps. The difference is that ours does not use skip connections between encoder and decoder, considering the potential streaming scenario [40]. Also, the bottleneck between the encoder and decoder that follows the UV layout of SMPL-X enables the face signal injection in our method. The runtime is compared in Table 3 where our method reaches a real-time framerate.
@@ 4.4 Ablation Study
Table 4: Choice of face signal. The 2D talking faces-based expression encoder from DPE [44] is a better choice as face signals.
@@
Choice of face signal. We conducted ablation experiments to validate the design choice of using a pretrained expression encoder from DPE [44] as the face signal. We compared to using a 3DMM as the face signal by extracting FLAME parameters from the frontal camera. We replaced the jaw pose of SMPL-X with that from the extracted FLAME parameters and used the expression parameters to condition our convolutional decoder. Note that in our method, the jaw pose is neutralized, i.e., the jaw movements are solely represented by the cVAE.
@@
The use of a pretrained encoder from 2D talking faces lets our method inherit the ability to both video-driven and audio-driven reenactment. With the help of SadTalker [74], we show the extension of DEGAS to full-body talking avatars. Given an audio clip, we firstly use SadTalker [74] to generate 2D talking videos from one face image of the subject, and then use the DPE [44] expression encoder to extract face signal to drive our avatars. The body pose generation is a whole another research track [69, 43, 9]. Here we only showcase the generation from TalkSHOW [69] in Figure 11. Noted that the jaw pose generated by TalkSHOW is not used.
@@ 5 Conclusion
In this paper, we proposed DEGAS, the first 3DGS-based method for full-body avatars with subtle and accurate facial expressions. We discussed in this paper that an expression latent space pretrained solely on 2D talking faces, is a better choice for the reenactment of 3D avatars, opening new possibilities for interactive life-like agents. The avatars modeled with DEGAS can be animated and rendered in real-time framerate to perform natural body motions and rich facial expressions. We conducted qualitative and quantitative experiments to validate the efficacy of our method. We also showed the audio-driven extension of our method to demonstrate its potential for downstream applications.
6 Acknowledgments
We thank the HKUST(GZ) DREAMS Lab for assisting in data collection. The work is supported by the National Social Science Foundation of China under grant number 24VWB020.
@@ C Ethics
We captured six human subjects in our proposed DREAMS-Avatar dataset. All subjects have given written consent for the captured footage in this dataset. We make the data publicly available for research purposes.